Sam acquires the following 5-year class property in 2018:

Asset Date of Acquisition Cost
X April 10 $ 60,000
Y July 25 $ 55,000
Z December 1 $ 110,000
$225,000

What is Sam's MACRS depreciation for 2018, assuming no Section 179 or bonus depreciation?
Answer #1

Since question isn't fully clear on the IRS convention, therefore, calculations are based on Mid-quarter convention with 200% declining balance method.

Depreciation on Property X = $60,000 * 25% = $15,000

Depreciation on Property Y = $55,000 * 15% = $8,250

Depreciation on Property Z = $110,000 * 0% = $0.